A tank contains 200 gallons of liquid. Five gallons of brine per minute flow into the tank, and each gallon of brine contains 2 pounds of salt. Five gallons of brine flow out of the tank per minute. Assume that the tank is kept well stirred.

A. Find a differential equation for the number of pounds of salt in the tank. Assuming the tank intially contains 50 pounds of salt, solve this differential equation.

B. How much salt is in the tank after one hour?

C. At what time will there be 399 lbs of salt in the tank?
Answer #1

5 Gallon/min J I 5 gallon 200 gallon liquid 50 pounds salt initially 2 found/gallon Let A (in pounds) be amount of salt at ti②. - put (=-350 ® in o the 400 -356 E uo &A= 400 - 350e to Pounds Answers 0 Put t=1 hour=bomins -60 DA(60) = 400 - 350 e = 32